OpenStudy (anonymous):

The sum of the interior angle measures of a convex polygon is 1440° How many sides does it have?

OpenStudy (anonymous):

is the answer 6 sides ?

OpenStudy (anonymous):

@ShadowLegendX @thomas5267 @zpupster @heavenfun13 @butterflydreamer

OpenStudy (aaronandyson):

(n-2)180 = 1440 n-2 = 8 n = 10 (# of sides)

OpenStudy (butterflydreamer):

The sum of the measures of the interior angle of a convex polygon with n sides is (n - 2 ) *180 . The angle sum they gave you was 1440 So (n - 2)*180 = 1440 Divide both sides by 180 So you get (n -2 ) = 8 then add 2 to both sides so n = 8+ 2 So n would be 10 :)